# Can Large Reasoning Models Self-Train?

### v1

@@ 4.1.1 What Happens After SRT Peaks in Performance?
After SRT’s performance peaks on the DAPO training dataset (see Figure 2), we observe that its test-time accuracy begins to deteriorate significantly. In fact, we find that a similar behavior also emerges when training on MATH-12k for more than two epochs, resulting in a noticeable performance collapse (see Appendix C for further details). To develop a clearer understanding of the underlying reasons for this phenomenon, in this section we investigate this model collapse closely.
First, we examine the training dynamics to ascertain whether the SRT self-training objective is being correctly optimized. We plot the running value of the SRT objective (Eqn. 6) in Figure 3. The observed performance collapse closely coincides with a sudden increase in the SRT self-reward objective, implying that the optimization procedure has, in fact, maximally optimized the training objective (self-consistency majority voting), despite a decline in actual output correctness. On the same figure, we further report the token-level average Kullback–Leibler divergence between the model under SRT training and the base model. We observe a sharp increase in KL divergence at the exact point when performance begins to deteriorate, indicating that the generative distribution of the model has substantially diverged from the original model.
These findings strongly suggest the occurrence of reward hacking—the model has learned to produce consistent responses in order to optimize its self-assigned reward, irrespective of their true correctness. Indeed, manual analysis of the model outputs (examples provided in Appnedix E) confirms this hypothesis: after collapse, the model outputs a very high entropy, essentially random, set of tokens followed by the same “template” final answer that is nearly independent of the input prompt. In other words, the initially strong correlation between the SRT objective and correctness is ultimately compromised, becoming no longer a reliable proxy signal.
Theoretical justification
There is a simple yet precise theoretical justification for this behavior:
The reinforcement learning optimization problem defined by the SRT objective
explicitly encourages consistency across outputs, independently of correctness.
Consequently, an optimal policy under this objective is to produce an identical response irrespective of the input, thus artificially attaining maximal possible reward. Therefore, one should naturally expect that continued training under such a proxy objective can result in this degenerate solution, especially when optimizing for this is easier than learning to solve the actual task.
4.2 Test-Time Self-Improvement
Figure 4: (Test-Time Self-Training Performance) Given the test dataset $\mathcal{D}_{\text{test}}$, one can perform SRT on $\mathcal{D}_{\text{test}}$ before making predictions. Our results show that this improves the majority voting performance on $\mathcal{D}_{\text{test}}$ without access to ground truth labels.
We compare the test-time performance of majority voting after SRT test-time training as well as without any test-time training. Empirically, we observe (Fig 4) that test-time training via SRT provides relatively limited, yet noticable, performance gains when measured under the maj@32 metric, compared to the popular majority voting baseline applied directly to outputs generated by the base model. Moreover, the performance gain over majority voting from the base model is larger on larger test datasets.
4.2.1 Why Doesn’t the Performance Collapse during Test-Time-Training?
@@
5 Can Model Collapse be Avoided?
As discussed in Section 4.1.1, the optimization objective in SRT can lead to significant initial improvements followed by eventual model collapse. Here, we explore complementary strategies to address model collapse and further enhance the performance achievable via self-training:
  1. (1)
  2. (2)
    An algorithmic strategy that mitigates the risk of model collapse by using pseudo-labels generated from a stable base model rather than from the continuously updated model.
  3. (3)
Additional results are discussed in Appendix D.
5.1 Early Stopping
Using a validation set for hyperparameter tuning and model selection is a well-established best practice. In our context, even a small labeled validation dataset can effectively identify the peak performance point during self-training, thereby mitigating model collapse. Figure 6 shows the progression of model performance, measured throughout training on the DAPO dataset and evaluated across several test sets. Crucially, we find that the peak performance consistently occurs around the same training step across different held-out datasets. This observation suggests that any of these datasets can serve effectively for early stopping. Concretely, the vertical line in Figure 6 illustrates early stopping based on just 1% of the DAPO dataset held out as a validation set. The model performance on all other evaluation datasets at this point remains close to optimal.
5.2 Self-Training with Offline-Generated Labels
Figure 7: (Self-Training with Offline-Generated Labels) Performance comparison between SRT and a baseline variant, where pseudo-labels are precomputed from a fixed base model checkpoint. The offline-generated labels maintain training stability while achieving comparable performance to SRT, highlighting a limitation of the online labeling strategy.
As identified in Section 4.1.1, the tendency toward model collapse arises because SRT prioritizes consistency over correctness, increasingly rewarding model agreement even if incorrect. A straightforward yet effective approach to address this involves generating pseudo-labels from a stable, previously fixed checkpoint, rather than leveraging labels from the evolving policy. To evaluate this approach, we generate pseudo-labels via majority voting rollouts from the Qwen2.5-Math-7B base model, store these offline-generated labels, and subsequently perform RL training against them. Figure 7 demonstrates that training with these offline-generated labels significantly stabilizes training while achieving performance comparable to SRT. Importantly, this indicates that the dynamic updating of pseudo-labels during training (online labeling) may not always confer substantial benefits, and instead can contribute to training instability.
5.3 Self-Training with Curriculum Learning
We hypothesize that model collapse occurs more rapidly when training on more challenging datasets. This conjecture aligns with our empirical findings in Section 4.1.1: the model experiences earlier collapse when training on the difficult DAPO dataset compared to the simpler MATH-12K dataset. The intuition is that, on a more challenging dataset, the model finds it easier to abandon its pretrained knowledge in favor of optimizing self-consistency rather than genuinely learning to solve the underlying task.
We leverage this hypothesis to implement a curriculum learning strategy (Bengio et al., 2009; Andrychowicz et al., 2017; Portelas et al., 2020; Florensa et al., 2017; Song et al., 2025; Lee et al., 2025; Tajwar et al., 2025) by identifying the ‘easiest’ subset of the DAPO dataset. To be precise, we retain 1/3-rd of the easiest DAPO prompts selected according to two distinct metrics:
  1. (1)
    Pass rate of the base model, which utilizes ground-truth labels.
  2. (2)
Figure 8 shows that training on these easier subsets significantly delays the onset of reward hacking, allowing for continuous improvement even across multiple epochs. Remarkably, performance on these curriculum subsets reaches levels comparable to standard RL training with ground-truth labels on the entire DAPO dataset. These promising results suggest that curriculum strategies may further extend the benefits of SRT, opening exciting avenues for future investigation.
@@ 6 Related Works
Previous works (Zelikman et al., 2022; Wang et al., 2023b; Huang et al., 2023; Madaan et al., 2023; Chen et al., 2024; Gulcehre et al., 2023; Singh et al., 2024; Ni et al., 2023; Hwang et al., 2024; Havrilla et al., 2024; Pang et al., 2024) have demonstrated the feasibility of LLMs’ self-improvement over their previous iteration by training on data distilled by the previous instances of the model. Most of these approaches usually have data filtering/reranking step in the pipeline, which is often performed by the model itself (Wu et al., 2025) or by training another (Hosseini et al., 2024) verifier model. Particularly, (Huang et al., 2023; Wang et al., 2023a; Prasad et al., 2024) demonstrated the feasibility of using majority voting and self-consistency to filter chain-of-thought traces that, when used as SFT training data, improve the LLM performance on downsteaming tasks. A concurrent work, (Zhao et al., 2025), proposes a self-evolution, self-play pipeline where an LLM generates coding problems of appropriate difficulty, solves and trains on them using RLVR. The model-generated solutions are validated by a code executor in the loop. To the best of our knowledge, our work is the first to explore online RLVR’s potential in self improving LLMs for math specific reasoning as opposed to a single round of SFT or DPO (Prasad et al., 2024; Huang et al., 2023). In a recent work, (Song et al., 2025) formalized a generation verification gap as central to the model’s ability to self-improve. Similarly, (Huang et al., 2025) proposed a “sharpening" mechanism as the key to self-improvement. Our SRT pipeline builds on top of both of these intuitions.
@@
Model collapse is a well-known phenomenon in training on self-generated training data (Alemohammad et al., 2024; Shumailov et al., 2024b, a; Bertrand et al., 2024; Briesch et al., 2025), and multiple approaches related to data mixing, reliable verification, training using contrastive loss using negative samples and curriculum learning have been proposed (Gerstgrasser et al., 2024; Feng et al., 2025; Briesch et al., 2025; Song et al., 2025; Gillman et al., 2024; Setlur et al., 2024) to prevent models from collapsing, which previous work on LLM’s easy to hard generalization (Lee et al., 2025) also utilize. However, in RL paradigm, we do not directly do supervised fine-tuning on model-generated data, and it remains an open question to what extent the previous findings of model collapse apply to our RLVR setting. In our work, we show that models trained using RL on self-labeled data often suffer from actor collapsing due to reward hacking (Amodei et al., 2016; Denison et al., 2024) and propose a few strategies to mitigate it.
@@
Several concurrent works look into the data efficiency of the RLVR pipeline. Notably, (Wang et al., 2025) shows that by just training on one example, the model can achieve performance equivalent to training on 1.2k examples from the DeepScaleR dataset (Luo et al., 2025). Similarly, TTRL (Zuo et al., 2025), also proposes a label-free online RLVR paradigm in a test-time setting, similar to SRT. Our work also explores this paradigm in Section 4.2, however, we focus on reducing the need for ground truth labels instead of training examples.
7 Conclusion
In this work, we introduced Self-Rewarded Training (SRT), a reinforcement learning approach enabling continual self-improvement of language models through intrinsic self-consistency, eliminating the need for external annotations. Our experiments demonstrated promising initial performance comparable to models trained with explicit ground-truth signals. However, we also identified significant limitations, notably reward hacking, where models optimize consistency rather than correctness, risking eventual performance collapse. We proposed preliminary mitigation strategies such as early stopping and curriculum learning to address this issue.
Further investigations could explore the role of the base model’s as well as develop more robust forms of self-verification. Additionally, integrating labeled data in a semi-supervised RL framework or exploring curriculum learning techniques are promising directions. We leave these as well as other exciting directions towards self-improving models for future work.
@@ Acknowledgements
This work has greatly benefited from the use of Delta’s advanced computing and data resource supported by the National Science Foundation (OAC 2005572) and the State of Illinois, as part of ACCESS-approved compute grants (Boerner et al., 2023). We also appreciate the computing resources of Bridges-2 (Brown et al., 2021) at Pittsburgh Supercomputing Center through ACCESS allocation CIS240901 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. Overall, this project used ACCESS grants CIS240901, CIS230366, CIS250216, and CIS250428 for its compute resources. Moreover, FT was partially supported by the U.S. Army Futures Command under Contract No. W519TC-23-C-0030 during the project.
The authors thank Brandon Pusateri, Jillian Lehosky, and Greg Bauer from ACCESS Support Staff for their incredible help at approving supplements and renewals for ACCESS compute grants throughout this project. Moreover, the work would not have finished so quickly without the help of Brett Bode from NCSA Delta Support Staff, who provided the authors with critical help in properly utilizing the Delta cluster. FT gratefully acknowledges Daman Arora, Yuda Song, Yiding Jiang, Ruiqi Zhang, Zhaoyi Zhou, Guanning Zeng, Yutong He and other members of the Zanette, Russ, and Auton lab for feedback and suggestions received on earlier versions of this work. FT also greatly benefited from using Prof. Aviral Kumar’s lab’s computational resources for initial implementation and debugging of the algorithm.
@@ A.1 Generation Verification Gap Through Majority Voting
Figure 9: Three of our test datasets display evidence of generation verification gap through majority voting. The positive gap between majority voting accuracy and per prompt accuracy means LLMs can utilize this gap as a learning signal to improve their average accuracy. x axis refers to the threshold cut off self-consistency (proportion of answers that are majority voted answers). Higher x value refers to more self-consistent LLM outputs. For example, at x = 0.4, we only keep LLM responses where at least 40% of the (properly parsed) responses were the majority voted answer. On y axis, average per prompt accuracy refers to the average number of correct answers out of the (successfully parsed) responses, across 32 generations per question (computed on a per prompt basis and then averaged over the dataset). Majority vote accuracy refers to how often the majority vote is the correct answer for that prompt (then averaged over the dataset). % of Prompts was computed over all 32 generations (not on the parsable answers for a fair comparison). If there is no prompt left over a certain threshold, the line plot ends there.
@@ Appendix B Details on Implementation and Training
B.1 RLOO Algorithm
All experiments in this work use RLOO(Ahmadian et al., 2024a; Kool et al., 2019) for its simplicity.
$\frac{1}{k}\sum_{i=1}^{k}[R(y_{(i)},x)-\frac{1}{k-1}\sum_{j\neq k}R(y_{(j)},x)]\nabla\log\pi(y_{(i)}|x)\text{ for }y_{(1)},...,y_{(k)}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\pi_{\theta}(\cdot|x)$
Here k is the number of online samples generated. RLOO creates a dynamic baseline for each sample without needing a separate value function, effectively estimating the expected return on-the-fly during training. Not having a value networks makes the training much simpler and (Ahmadian et al., 2024a) shows it to be a very strong baseline. For all experiments in this paper, we set k=32.
In our implementation, we did not add KL penalty to the loss function, rather to the reward itself. In verl framework, this can be configured using algorithm.use_kl_in_reward=True and actor_rollout_ref.actor.use_kl_loss=False. To estimate KL penalty, we use the low variance KL estimator proposed by Schulman (2020):

### v2

@@ 4.2 Can Self-Improvement from SRT be Sustained Indefinitely?
Figure 8: (Sweep over training configurations) We vary the training settings for SRT from the default ones described in Appendix B.3 and record the observations. All experiments are run using Qwen2.5-Math-7B on DAPO. Evaluations are done using AIME 2025, but other datasets show similar behavior. In summary, choice of RL algorithm between GRPO and RLOO, and different KL coefficients do not affect model collapse significantly. Lowering learning rate slows down model collapse as expected. Finally, reducing number of generations injects noise into the majority vote, surprisingly resulting in a slow down of model collapse.
Additional experiments and ablations.
We have run additional experiments and ablations to study the behavior of self-training under different training scenarios. Figure 8 shows our main findings: (1) choice of RL algorithm (GRPO vs RLOO) does not affect the final outcome of SRT-training, (2) increasing KL coefficient to incentivize the model to stay close to the base policy also does not mitigate reward hacking, as the training signal from the reward hacked solution is too strong, (3) decreasing learning rate seems to delay model collapse but not eliminate it, and we hypothesize that prolonged training with lower learning rates would still result in complete model collapse, and (4) surprisingly, reducing the number of generations per prompt injects noise into the training signal, which delays quick model collapse by exploiting the majority voting answer. Additional experiments related to SRT training, including training curves on all test datasets and effect of tuning additional hyperparameters like entropy coefficient, can be found in Appendix C and G. We also investigated various mitigation strategies for SRT-induced model collapse, including early stopping, curriculum learning, and employing a fixed teacher to generate labels – refer to Appendix I for our findings.
Takeaway 2: Self-Training benefits may not extend indefinitely The question of whether self-training can be extended indefinitely has mixed results: while under controllable difficulty, SRT can keep improving beyond the base model on progressively more difficult tasks, training on real-world math problems demonstrate the phenomenon of reward hacking — sustained self-improvement requires developing additional regularization measures to be effective.
@@ 5 Related Works
Previous works (Zelikman et al., 2022; Wang et al., 2023b; Huang et al., 2023; Madaan et al., 2023; Chen et al., 2024; Gulcehre et al., 2023; Singh et al., 2024; Ni et al., 2023; Hwang et al., 2024; Havrilla et al., 2024; Pang et al., 2024a) have demonstrated the feasibility of LLMs’ self-improvement over their previous iteration by training on data distilled by the previous instances of the model. Most of these approaches usually have data filtering/reranking step in the pipeline, which is often performed by the model itself (Wu et al., 2025b) or by training another (Hosseini et al., 2024) verifier model. Particularly, (Huang et al., 2023; Wang et al., 2023a; Prasad et al., 2024) demonstrated the feasibility of using majority voting and self-consistency to filter chain-of-thought traces that, when used as SFT training data, improve the LLM performance on downsteaming tasks. A concurrent work, (Zhao et al., 2025a), proposes a self-evolution, self-play pipeline where an LLM generates coding problems of appropriate difficulty, solves and trains on them using RLVR. The model-generated solutions are validated by a code executor in the loop. Previous works have studied self-improvement by generating labels through majority voting  (Prasad et al., 2024; Huang et al., 2023), but these works are typically confined to one or a few rounds of SFT or DPO (Rafailov et al., 2024). This essentially involves distilling majority voting labels from a fixed policy into the current policy over each round of training. In contrast, we explore online RLVR’s potential in self-improving LLMs where the label generating policy evolves after every gradient step. A few concurrent works (Chen et al., 2025; Prabhudesai et al., 2025; Zhao et al., 2025b; Shao et al., 2025) have also explored various forms of self-rewarding mechanisms through majority voting or a similar metric of self-consistency (e.g., token/sequence level entropy) in Online RLVR. Finally, in a recent work, (Song et al., 2025b) formalized a generation verification gap as central to the model’s ability to self-improve. Similarly, (Huang et al., 2025) proposed a “sharpening" mechanism as the key to self-improvement. Our SRT pipeline builds on top of both of these intuitions. We refer the interested reader to Gao et al. (2025) for a survey of other works associated with self-evolving LLM agents.
@@
Model collapse is a well-known phenomenon in training on self-generated training data (Alemohammad et al., 2024; Shumailov et al., 2024b; a; Bertrand et al., 2024; Briesch et al., 2025), and multiple approaches related to data mixing, reliable verification, training using contrastive loss using negative samples and curriculum learning have been proposed (Gerstgrasser et al., 2024; Feng et al., 2025; Briesch et al., 2025; Song et al., 2025b; Gillman et al., 2024; Setlur et al., 2024) to prevent models from collapsing, which previous work on LLM’s easy to hard generalization (Lee et al., 2025) also utilize. However, in RL paradigm, we do not directly do supervised fine-tuning on model-generated data, and it remains an open question to what extent the previous findings of model collapse apply to our RLVR setting. In our work, we show that models trained using RL on self-labeled data often suffer from actor collapsing due to reward hacking (Amodei et al., 2016; Denison et al., 2024) and propose a few strategies to mitigate it. Finally, a concurrent and complementary work, Zhang et al. (2025b), has also demonstrated the failure of various self-reward mechanisms to sustain self-improvement under prolonged training, which further validates the observations in this work.
@@
Several concurrent works look into the data efficiency of the RLVR pipeline. Notably, (Wang et al., 2025) shows that by just training on one example, the model can achieve performance equivalent to training on 1.2k examples from the DeepScaleR dataset (Luo et al., 2025). Similarly, TTRL (Zuo et al., 2025), also proposes a label-free online RLVR paradigm in a test-time setting, similar to SRT. Our work also explores this paradigm in Section H.1 as an extension of SRT.
6 Limitations and Conclusion
In this work, we examine a simple strategy of leveraging an LLM’s self-consistency to train it via an RL framework. Our experiments on synthetic reasoning tasks with controllable difficulty levels demonstrate the promise of such a framework: not only can LLMs improve their performance on these tasks, but they can also improve the quality of their self-supervision for the next training step. Our analysis, however, also reveals the limitation of leveraging self-consistency as training reward: prolonged training under such a framework can lead to reward hacking and complete model collapse. Future investigations could explore how to develop more robust forms of verification, extend self-verification to other domains like coding, and curriculum learning strategies that expose the model to problems of only the right difficulty. Additionally, leveraging LLM-as-judges or generative verifiers to improve the training signal, or employing additional consistency regularization between the chain-of-thought and the final answer to mitigate reward hacking can be promising future research. We leave these and other promising directions for sustained self-improvement for future work.
@@ Acknowledgements
This work has greatly benefited from the use of Delta’s advanced computing and data resource supported by the National Science Foundation (OAC 2005572) and the State of Illinois, as part of ACCESS-approved compute grants (Boerner et al., 2023). We also appreciate the computing resources of Bridges-2 (Brown et al., 2021) at Pittsburgh Supercomputing Center through ACCESS allocation CIS240901 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. Overall, this project used ACCESS grants CIS240901, CIS230366, CIS250216, and CIS250428 for its compute resources. Moreover, FT was partially supported by the U.S. Army Futures Command under Contract No. W519TC-23-C-0030 during the project. Andrea Zanette was partially supported by the National Science Foundation under Grants CCF-2106778 and DMS-2134080.
The authors thank Brandon Pusateri, Jillian Lehosky, and Greg Bauer from ACCESS Support Staff for their incredible help at approving supplements and renewals for ACCESS compute grants throughout this project. Moreover, the work would not have finished so quickly without the help of Brett Bode from NCSA Delta Support Staff, who provided the authors with critical help in properly utilizing the Delta cluster. FT gratefully acknowledges Daman Arora, Yuda Song, Yiding Jiang, Ruiqi Zhang, Zhaoyi Zhou, Guanning Zeng, Yutong He and other members of the Zanette, Russ, Auton, and AIRe lab for feedback and suggestions received on earlier versions of this work.
@@ A.1 Generation Verification Gap Through Majority Voting
Figure 9: Three of our test datasets display evidence of generation verification gap through majority voting. The positive gap between majority voting accuracy and per prompt accuracy means LLMs can utilize this gap as a learning signal to improve their average accuracy. x axis refers to the threshold cut off for self-consistency (proportion of answers that are majority voted answers). Higher x value refers to more self-consistent LLM outputs. For example, at x = 0.4, we only keep LLM responses where at least 40% of the (properly parsed) responses were the majority voted answer. On y axis, average per prompt accuracy refers to the average number of correct answers out of the (successfully parsed) responses, across 32 generations per question (computed on a per prompt basis and then averaged over the dataset). Majority vote accuracy refers to how often the majority vote is the correct answer for that prompt (then averaged over the dataset). % of Prompts was computed over all 32 generations (not on the parsable answers for a fair comparison). If there is no prompt left over a certain threshold, the line plot ends there.
@@ Appendix B Details on Implementation and Training
B.1 RL Algorithms
We use two RL algorithms in this work: RLOO(Ahmadian et al., 2024; Kool et al., 2019) and GRPO (Shao et al., 2024). Recent work (Oertell et al., 2024) has shown that heuristic policy gradient algorithms like GRPO can produce unexpected results by increasing or decreasing reasoning performance even under random rewards, where policy gradient should be zero in expectation, and that RLOO does not have this problem. Since SRT is compatible with both RL algorithms, we experiment with both and observe no noticeable difference in the resulting behavior.
In our implementation (verl), we use the following RL objective for both RLOO and GRPO:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\ \{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Big($ $w_{i,t}(\theta)\hat{A}_{i,t},$
$\text{clip}(w_{i,t}(\theta),\ 1-\varepsilon,\ 1+\varepsilon)\hat{A}_{i,t}\Big)\Bigg]$
where $w_{i,t}(\theta)$ is the importance ratio, defined as:
$w_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid x,\ y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x,\ y_{i,<t})}$
Since we operate fully on-policy, i.e., one RL step per one batch of generated rollouts, this is always one in our experiments. The same advantage defined at a sequence level is applied to each token in the sequence, so henceforth we will drop the $t$ from the notation as well.
The main difference between GRPO and RLOO then stems from their use of different advantage functions. RLOO objective uses the following advantage function:
$\frac{1}{G}\sum_{i=1}^{G}[R(y_{(i)},x)-\frac{1}{G-1}\sum_{j\neq k}R(y_{(j)},x)]$
whereas GRPO uses the following advantage function:
$\hat{A}_{i}=\frac{r(x,y_{i})-\text{mean}\left(\{r(x,y_{i})\}_{i=1}^{G}\right)}{\text{std}\left(\{r(x,y_{i})\}_{i=1}^{G}\right)}$
Here $G$ is the number of online samples generated. Both RLOO and GRPO create a dynamic baseline for each sample without needing a separate value function, effectively estimating the expected return on-the-fly during training. Not using a separately trained value network (as opposed to PPO) makes the training much simpler for both algorithms.
In our implementation, we did not add the KL penalty to the loss function, but rather to the reward itself while running RLOO, following recent work such as Tang & Munos (2025). In verl framework, this can be configured using algorithm.use_kl_in_reward=True and actor_rollout_ref.actor.use_kl_loss=False. However, this does not work for GRPO due to advantage normalization by the standard deviation, and so for GRPO we add KL penalty to the loss function directly. To estimate KL penalty, we use the low variance KL estimator proposed by Schulman (2020):